# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

### v3

@@ Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. While considerable success has been achieved, there still exist urgent limitations. Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes. Our codes are available at https://github.com/itsnotacie/TOMM2024-COMAE.
Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. While considerable success has been achieved, there still exist urgent limitations. Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes. Our source codes are available at https://github.com/YihangZhou0424/NN2024-COMAE.
@@
Prototype Learning, Local Representation, Zero-shot Hashing, Image Zero-shot Hashing, Image Hashing, Image Retrieval.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXjournal: TOMMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Image search
@@
Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Lew et al., 2006; Wang et al., 2017; Ju et al., 2024). Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Ju et al., 2023). The capability to depict intricate data in a compact binary format not only accelerates retrieval processes but also markedly improves storage efficiency. Consequently, image hashing emerges as a highly valuable asset in situations that necessitate swift and resource-efficient information retrieval.
@@
With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017; Long et al., 2021). Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue. Visual attributes, pervasive across prevalent image datasets, serve to delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022). Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks (Arik and Pfister, 2019; Xu et al., 2017). Fig. 1 shows an example of the class-level attributes. Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager. AH (Xu et al., 2017) adopts a straightforward approach by constructing a multi-task architecture to capture the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
@@
While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exists urgent limitations (Li and van Gemert, 2021; Xu et al., 2017). Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021), as the locality exhibits more effective transferability between seen classes and unseen classes. However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms. Secondly, existing works mainly focus on individual image representation learning, optimizing each instance independently. Each dimension of the visual attribute is a continuous value. In contrast to the binary one-hot class encoding, these continuous values encompass richer and more nuanced distance-related information. These informative attributes are aptly poised to capture intricate pairwise relationships, while they are largely ignored by prevailing works. Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels. Yet their actual relationships are more complicated than such simple presumes. To be specific, two images may belong to the same class but pose disparate attributes. For example, two images with a “bird” class may have “black leg” and “yellow leg”. Similarly, instances with different classes may also exhibit sharing analogous attributes. For example, two images with “bird” and “wapiti” classes may have the same “black leg”. The existing methods may not lead to the global optimum, thus necessitating performing joint optimization to achieve the overall optimum for both tasks.
@@
To tackle the above issues, we conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE. It combines relationships among visual attributes, image features, class labels, and binary hash codes. Specifically, we develop a supervised learning objective that localizes visual attributes for point-wise consistency. By utilizing the attribute prototype network, we aim to capture the fine-grained details and distinctive characteristics of images, enabling more accurate and informative hash codes. To comprehensively harness numerical attribute features, we construct positive and negative instances for each image based on the numerical disparities of attributes and utilize contrastive learning to further model their distance relationships. Thus, our model can capture the subtle variations in attribute values across different images and learn representations that encode the common visual characteristics associated with these attributes. Finally, class-wise optimization is proposed to cohesively learn the hash code, image representation, and visual attributes more effectively, which fosters improved generalization by considering the relationships and dependencies between different classes during the learning process. We jointly learn discriminative global and local features based on the class-level attributes. Extensive experiments on the ZSH benchmark datasets show that COMAE outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes. It serves to validate the effectiveness of comprehensive attribute explorations.
Figure 1. An example of the class-level attributes.
Our contributions are summarized as follows:
@@
  • We conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE, it depicts the relationships between seen classes and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints.
@@
  • We design an attribute prototype network, a contrastive learning unit, and a joint-optimal label utilization to implement the above three components.
@@
This paper is generally organized as follows. We begin with briefly reviewing related works in Section 2. Then the proposed COMAE is technically introduced in Section 3. We provide extensive experiment results and discussions in Section 4. Finally, we conclude this work in Section 5.
@@ 1.1. Deep Hashing
Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000). Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008) and ITQ (Gong et al., 2012). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
Over the past decades, deep learning models have shown excellent progress in the field of Computer Vision (Khan et al., 2018; Pourpanah et al., 2022) and Natural Language Processing (Song et al., 2020; Long et al., 2021), many scholars pay attention to leveraging deep neural networks to learn the abstract representations. Deep hashing methods can be achieved through three paths. The first path utilizes limited information, which has a lower upper bound of the capability. ASZH (Shi et al., 2022) is a typical work in the first path, that learns hash functions based on the seeable category labels, for generating binary codes for unseeable data. The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseeable classes. TSK, SitNet, and OPZH (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space. However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough. The third path (Xu et al., 2017) constructs relationships among attributes, image features, binary codes, and class labels. These models are considered to be more generalized when dealing with out-of-distribution and open-set samples (Arik and Pfister, 2019; Gao et al., 2019).
@@ 1.2. Zero-shot Learning
Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022; Long et al., 2020) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers. Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022) harvest global visual features from networks, whether pre-trained or capable of end-to-end training. Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications. Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020), as they struggle to discern the subtle variances between seen and unseen categories. Recent innovations in ZSL predominantly concentrate on methods based on attention (Zheng and Gupta, 2022; Sylvain et al., 2019; Xu et al., 2020), employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird. This emphasis highlights the critical roles of specificity and the arrangement of image features in ZSL. Proposals for prototype networks (Sylvain et al., 2019) strive to pinpoint various attributes within images to bolster the precision of feature localization. Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes (Xu et al., 2020).
1.3. Contrastive Learning
Contrastive learning aims at honing representations by juxtaposing similar instance pairs (positive pairs) against dissimilar ones (negative pairs) (Hadsell et al., 2006; Han et al., 2021), enhancing the model’s ability to differentiate between closely related and distinct instances effectively. Many works demonstrate that contrastive learning is able to alleviate the problems of numerous categories and imbalanced class distributions. In addition, it can amplify the subtle distinctions among similar classes, which assists in capturing intricate details that are often overlooked by basic models. (Hadsell et al., 2006) is the first attempt work to improve learned representations by contrasting positive and negative pairs. SimCLR (Chen et al., 2020) utilizes a self-supervised learning architecture that substitutes the conventional memory bank with batch-specific elements, significantly enhancing its performance on the ImageNet dataset. As hash code is a form of representation, recent researches (Yu et al., 2022; Song et al., 2023; Qiu et al., 2021) take attempts to utilize contrastive learning in deep hashing. (Yu et al., 2022) proposes weighted contrastive hashing to explore the fine-grained semantic relations among images. (Song et al., 2023) propose to However, random sampling could result in augmented images with different semantics, deteriorating the performance of contrastive learning.
2. Method
The architecture of the proposed COMAE is shown in Fig. 2. Our proposed COMAE depicts the relationships between seen and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints. To improve the image locality and attribute representations, we propose a point-wise objective. To learn the representations from individual learning to context-based learning, we utilize contrastive learning to build pair-wise loss. For capturing relationships between attributes and class labels, a class-wise constraint is designed.
@@ 2.1. Problem Definition
Let $X=\{x_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times H\times W\times C}$ be $N$ images, where $H,W,C$ denote the image height, width, and channel. The matrix $Y=\{y_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ is the label matrix, where $C$ denotes the number of classes. The classes are divided into the seeable class $Y^{s}$ and the unseeable class $Y^{u}$, where $|Y^{s}|\cup|Y^{u}|=|C|,|Y^{s}|\cap|Y^{u}|=\emptyset$. Let $A=\{a_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times K}$ denote the class-level attribute vector in $K$ dimensions. By using training samples which contain labeled images and attributes from seeable classes, i.e., $S=\{x,y,a|x\in X,y\in Y^{s},a\in A\}$, we aim to learn a hashing model $\mathcal{H}\colon X\rightarrow b_{i}\in\{-1,1\}^{l}$, where $b_{i}$ represents the learned $l$-bit binary hash codes. In the testing process, attribute vectors of unseen classes are also known. In this paper, zero-shot hashing focuses on predicting the hashing label of images from unseen classes, i.e., $X\rightarrow Y^{u}$.
@@ 2.2. Point-wise Attribute Consistency
Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021). For more subtly depicted image and attribute characterizations, we first conduct point-wise consistency constraint, which takes the attribute regression loss as the supervision. Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples (Xu et al., 2020; Tan et al., 2022) in recent years. In the prototype learning scheme, each category is represented by a prototype. This prototype can be the mean of the sample in the category, the center of mass, or some other representative sample. Thus we introduce a prototype module aiming at capturing the relationships of local features and visual attributes.
Let $f_{i,j}(x)\in\mathbb{R}^{C}$ denote the image representation at spatial position (or location) $(i,j)$. It depicts the information on local image regions. Then we learn sorts of attribute prototypes $P=\{p_{k}\in\mathbb{R}^{C}\}_{k=1}^{K}$ to predict attributes from the local visual characterizations. Let $p_{k}$ denote the learnable prototype embeddings for the $k$-th attribute. As shown in Fig. 2, $p_{1}$ and $p_{2}$ correspond to the prototypes for “red beak” and “yellow leg” respectively. For each attribute, we produce a similarity map $M^{k}\in\mathbb{R}^{H\times W}$, where each element is computed by a non-linear network $\Theta(\cdot)$ between for $p_{k}$ and each local feature. To associate each visual attribute with its closest local feature, we gain the predicted $k$-th attribute $\hat{a_{k}}$ by taking the maximum value in the similarity map $M^{k}$,
@@
Furthermore, we refine the similarity map to intensively focus on its peak region. To achieve this, we employ a compactness regularizer (Zheng et al., 2017) on each similarity map $M^{k}$, ensuring a concentrated similarity assessment. The equation is defined as,
(3) $\mathcal{L}_{\text{reg}}=\sum^{K}_{k}\sum^{H}_{h}\sum^{W}_{w}M_{i,j}^{k}[(i-\tilde{i})^{2}+(j-\tilde{j})^{2}],$
where $(\tilde{i},\tilde{j})=\arg\max_{i,j}M_{i,j}^{k}$ identifies the coordinate with the highest value in $M^{k}$, which ensuring the attribute prototype mirrors only a minimal subset of local features, thus achieving a more streamlined similarity map. This method emphasizes the precision in the prototype’s focus, leading to a denser representation of similarity. Then the $\mathcal{L}_{\text{pointwise}}$ is changed to $\mathcal{L}_{\text{pointwise}}=\sum\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}+\mathcal{L}_{\text{reg}}$. The effective representations of image locality ensure a more robust and versatile transfer of knowledge, contributing to the overall effectiveness of COMAE.

### v5

@@ Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. However, existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Furthermore, a theoretical analysis is provided to show the effectiveness. Experimental results demonstrate that COMAE outperforms state-of-the-art hashing models, especially in scenarios with a larger number of unseen label classes. Our codes are available at https://github.com/itsnotacie/ICMR2025-COMAE.
@@
Deep Hash, Zero-shot Hashing, Attribute Learning, Prototype Learning, Image Retrieval, Image Hashing
journalyear: 2025copyright: ccconference: Proceedings of the 2025 International Conference on Multimedia Retrieval; June 30-July 3, 2025; Chicago, IL, USAbooktitle: Proceedings of the 2025 International Conference on Multimedia Retrieval (ICMR ’25), June 30-July 3, 2025, Chicago, IL, USAdoi: 10.1145/3731715.3733294isbn: 979-8-4007-1877-9/2025/06ccs: Information Systems Similarity Measuresccs: Information Systems Information retrievalccs: Information systems Image searchccs: Computing methodologies Visual content-based indexing and retrieval
1. Introduction
@@
Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Wang et al., 2017; Dong et al., 2024a; Arik and Pfister, 2019; Luo et al., 2025; Long et al., 2021; Li et al., 2025c, d; Zeng et al., 2025). Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019; Long et al., 2020).
@@
With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017). Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue. Visual attributes, pervasive across prevalent image datasets, delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022; Ma et al., 2024). Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks (Arik and Pfister, 2019; Xu et al., 2017). Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager. AH (Xu et al., 2017) adopts a straightforward approach by constructing a multi-task architecture to capture the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
@@
While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exist urgent limitations (Li and van Gemert, 2021; Xu et al., 2017). Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021a), as the locality exhibits more effective transferability between seen classes and unseen classes. However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms. Secondly, existing works mainly focus on individual image representation learning, optimizing each instance independently. Each dimension of the visual attribute is a continuous value. In contrast to the binary one-hot class encoding, these continuous values encompass richer and more nuanced distance-related information. These informative attributes are aptly poised to capture intricate pairwise relationships, while they are largely ignored by prevailing works. Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels. Yet their actual relationships are more complicated than such simple presumptions. To be specific, two images may belong to the same class but pose disparate attributes. For example, as shown in Fig. 1, two images with a “bird” class may have “black leg” and “brown leg”. Similarly, instances with different classes may also exhibit analogous attributes. For example, two images with “bird” and “wapiti” classes may have the same “black leg”. The existing methods may not lead to the global optimum, thus necessitating performing joint optimization to achieve the overall optimum for both tasks.
@@
To tackle the above issues, we conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE. It combines relationships among visual attributes, image features, class labels, and binary hash codes. Specifically, we develop a supervised learning objective that localizes visual attributes for point-wise consistency. By utilizing the attribute prototype network, we aim to capture the fine-grained details and distinctive characteristics of images, enabling more accurate and informative hash codes. To comprehensively harness numerical attribute features, we construct positive and negative instances for each image based on the numerical disparities of attributes and utilize contrastive learning to further model their distance relationships. Thus, our model can capture the subtle variations in attribute values across different images and learn representations that encode the common visual characteristics associated with these attributes. Finally, class-wise optimization is proposed to cohesively learn the hash code, image representation, and visual attributes more effectively, which fosters improved generalization by considering the relationships and dependencies between different classes during the learning process. We jointly learn discriminative global and local features based on the class-level attributes. Extensive experiments on the ZSH benchmark datasets show that COMAE outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes. Our contributions are summarized as follows:
@@
  • We conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE, which depicts the relationships between seen classes and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints.
@@
  • We design an attribute prototype network, a contrastive learning unit, and a joint-optimal label utilization to implement the above three components, with theoretical analysis to show its effectiveness.
@@ 2.1. Deep Hashing
Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000; Dong et al., 2024b; Qin et al., 2024; Chen et al., 2025a). Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008) and ITQ (Gong et al., 2012). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
Over the past decades, deep learning models have shown excellent progress in the field of computer vision (Zhu et al., 2024a, b; Dong et al., 2024a; Chu et al., 2025; Qin et al., 2023), language processing (Wu et al., 2024; Guo et al., 2025; Cai et al., 2023; Xiao et al., 2025; Lu et al., 2024), and multimodal tasks (Lin et al., 2024; Li et al., 2021b; Xiao et al., 2021; Li et al., 2024, 2025b). Deep hashing methods can be achieved through three paths. The first path utilizes limited information, which has a lower upper bound of the capability. ASZH (Shi et al., 2022; Qin et al., 2022) is a typical work in the first path, which learns hash functions based on the visible category labels, for generating binary codes for unseeable data. The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseen classes. TSK, SitNet, OPZH, and PIXEL (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019; Dong et al., 2024a) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space. However, PIXEL introduced BERT, which resulted in longer processing times. However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough. The third path (Xu et al., 2017) constructs relationships among attributes, image features, binary codes, and class labels. These models are considered to be more generalized when dealing with out-of-distribution and open-set samples (Arik and Pfister, 2019; Gao et al., 2019).
@@ 2.2. Zero-shot Learning
Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers. Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022; Chen et al., 2025b) harvest global visual features from networks, whether pre-trained or capable of end-to-end training. Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications. Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020; Li et al., 2025a), as they struggle to discern the subtle variances between seen and unseen categories. Recent innovations in ZSL predominantly concentrate on methods based on attention (Zheng and Gupta, 2022; Sylvain et al., 2019; Xu et al., 2020), employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird. This emphasis highlights the critical roles of specificity and the arrangement of image features in ZSL. Proposals for prototype networks (Sylvain et al., 2019) strive to pinpoint various attributes within images to bolster the precision of feature localization. Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes (Xu et al., 2020).
3. Model: COMAE
The architecture of the proposed COMAE is shown in Fig. 1. Our proposed COMAE depicts the relationships between seen and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints. To improve the image locality and attribute representations, we propose a point-wise objective. To learn the representations from individual learning to context-based learning, we utilize contrastive learning to build a pair-wise loss. For capturing relationships between attributes and class labels, a class-wise constraint is designed.
@@ 3.1. Problem Definition
Let $X=\{x_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times H\times W\times C}$ be $N$ images, where $H,W,C$ denote the image height, width and channel. The matrix $Y=\{y_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ denotes the labels, where $C$ denotes the number of classes. The labels are divided into the seeable class $Y^{s}$ and the unseeable class $Y^{u}$, where $|Y^{s}|\cup|Y^{u}|=|C|,|Y^{s}|\cap|Y^{u}|=\emptyset$. Let $A=\{a_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times K}$ denote the class-level attribute vector in $K$ dimensions. By using training samples which contain labeled images and attributes from seeable classes, i.e., $S=\{x,y,a|x\in X,y\in Y^{s},a\in A\}$, we aim to learn a hashing model $\mathcal{H}\colon X\rightarrow b_{i}\in\{-1,1\}^{l}$, where $b_{i}$ denotes learned $l$-bit binary hash codes. For the testing process, attribute vectors of unseen classes are also known. In this paper, zero-shot hashing focuses on predicting the hashing label of images from unseen classes, i.e., $X\rightarrow Y^{u}$.
@@ 3.2. Point-wise Attribute Consistency
Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021a). For more subtly depicted image and attribute characterizations, we first conduct a point-wise consistency constraint, which takes the attribute regression loss as the supervision. Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples (Xu et al., 2020; Tan et al., 2022) in recent years. In the prototype learning scheme, each category is represented by a prototype. This prototype can be the mean of the sample in the category, the center of mass, or some other representative sample. Thus we introduce a prototype module aiming at capturing the relationships of local features and visual attributes.
Let $f_{i,j}(x)\in\mathbb{R}^{C}$ denote the image representation at spatial position (or location) $(i,j)$. It depicts the information on local image regions. Then we learn sorts of attribute prototypes $P=\{p_{k}\in\mathbb{R}^{C}\}_{k=1}^{K}$ to predict attributes from the local visual characterizations. Let $p_{k}$ denote the learnable prototype embeddings for the $k$-th attribute. As shown in Fig. 1, $p_{1}$ and $p_{2}$ correspond to the prototypes for “orange crown” and “yellow skin” respectively. For each attribute, we produce a similarity map $M^{k}\in\mathbb{R}^{H\times W}$, where each element is computed by a non-linear network $\Theta(\cdot)$ between $p_{k}$ and each local feature. To associate each visual attribute with its closest local feature, we gain the predicted $k$-th attribute $\hat{a_{k}}$ by taking the maximum value in the similarity map $M^{k}$,